OpenStudy (anonymous):

My name is John. I have never been here before. I am having trouble understanding a sample problem provided in my text. It explains calculating the Rate in a simple interest problem.

OpenStudy (anonymous):

I attached an image of my issue

OpenStudy (anonymous):

my first issue is... isn't the formula upside down? my second issue is... where are the decimal places coming from?

OpenStudy (anonymous):

I thought it was Interest over Principal X Time And I thought 1750 X 5 was 8,750

OpenStudy (anonymous):

ok, first of all, you should understand how the simple interest formula works. it goes like this: I=P*R*T when I is the interest you get after you inverst amount P at a rate R (rate can be found by dividing the percentage by 100) for time T in years. so..

OpenStudy (anonymous):

you can rearange the formula, to find rate like this: I=P*R*T R=I/(PT)

OpenStudy (anonymous):

because you divide both sides of I=PRT by PT

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

I get it but it reads R=(PT)/I

OpenStudy (anonymous):

or am i just not inturpriting it right?

OpenStudy (anonymous):

the formula is wrong in your example it should by R=I/(PT)

OpenStudy (anonymous):

and what about the 1750 X 5 = 87.50?

OpenStudy (anonymous):

the rest is correct, it is just that the formula is wrong, but the final answer is correct

OpenStudy (anonymous):

i realize the final answer was correct by using an online calculator :) AND I think they worked it out using the correct formula. I just don't understand some if there math

OpenStudy (anonymous):

I don't understand how 1750 * 2 1/2 = 87.5

OpenStudy (anonymous):

oops, i ment 3500 * 2 1/2 = 87.5

OpenStudy (anonymous):

well, the way these guys wrote it in that paper is completely nonsensical, so just delete it because the method of solving there is wrong and confusing

OpenStudy (anonymous):

you should get some trustworthy questions, not stuff like that

OpenStudy (amistre64):

simple interest is calculated at: I = Prt

OpenStudy (anonymous):

Ok, thats what I needed to here :) This was provided to me by an online government career development program. Where is your "Thumbs up" button?

OpenStudy (anonymous):

amistre, I already covered that abouve

OpenStudy (amistre64):

...but..but... it so far up there......

OpenStudy (anonymous):

it is near my name

OpenStudy (anonymous):

=D LOL

OpenStudy (anonymous):

by chance do you have a short explanation why this formula assumes 360 days in a year instead of 365?

OpenStudy (anonymous):

it does not

OpenStudy (amistre64):

accountants cant count? :)

OpenStudy (anonymous):

i used the word formula for lack of better term maybe i should have said concept

OpenStudy (amistre64):

it might have something to do with the Jewish culture, they started alot of this interest bearing stuff and they had 360 day years...

OpenStudy (anonymous):

but the simple interest formula assumes 365, I am certain

OpenStudy (anonymous):

they have provided me with more nonsense I assume :)

OpenStudy (anonymous):

sure they have, that example you showed was enough to prove that those guys are MORONS, who ever they are

OpenStudy (anonymous):

they translated 36 days as 36/360 for this formula

OpenStudy (anonymous):

if you are interested in learning maths fast, and in a fun way, I know a great site

OpenStudy (anonymous):

im interested in relearning math. I quit when I got to calculus years ago because I had poor trig skills

OpenStudy (anonymous):

http://www.khanacademy.org this is the site. IT IS 100% free, I LOVE IT. It covers everything from calculus to percentages to biology

OpenStudy (anonymous):

WATCH the intro vid

OpenStudy (anonymous):

you will get an idea how to use it

OpenStudy (anonymous):

http://www.khanacademy.org/#browse

OpenStudy (anonymous):

here is the list of topics, when you click on topics, you are directed to videos

OpenStudy (anonymous):

and above, the is a button practics, if you click on it you will be given some questions

OpenStudy (anonymous):

I love that its free. I paid $600 for this nonsense schooling :P

OpenStudy (anonymous):

well, just watch the intro video at the homepage of http://www.khanacademy.org/#browse

OpenStudy (amistre64):

Simple interest normally uses a 360-day year for simplicity, i.e., 30 day months. Interest calculated this way earns a little more than if using a 365-day year. Here is an example for a 90 day loan ... from google

OpenStudy (anonymous):

yea, but it depends on a problem and scenario, believe me, my text book uses 365

OpenStudy (anonymous):

I read banks use 360 and the federal reserve uses 365

OpenStudy (anonymous):

http://www.khanacademy.org/ this video intro is like 20 min, but it is cool, you will understand how the site works, it also has bill gate at the end :)

OpenStudy (anonymous):

So, again, where are these thumbs up fan buttons?

OpenStudy (amistre64):

just click"become a fan" next to the screen name its in blue

OpenStudy (anonymous):

it is the blue "Become a fan" button

OpenStudy (anonymous):

do you see it?

OpenStudy (anonymous):

Im gussing my adblocker is hiding it

OpenStudy (anonymous):

thats because you only have one post/question created, if you ask another question, the fan system should be unlocked for you

OpenStudy (anonymous):

but it is not a big deal anyways =D just a game :)

OpenStudy (anonymous):

is it inline on this page, or supposed to be in the info that appears on hover?

OpenStudy (anonymous):

Oh ok, odd says im a lifesaver?

OpenStudy (anonymous):

inline, and on the info, but as i have said, you have to have atleast a few questions. You are a lifesaver because I faned you :)

OpenStudy (anonymous):

today, yours should say sanity saver

OpenStudy (anonymous):

well, just use the http://www.khanacademy.org site to learn stuff, using those videos I learned calculus derivatives like in 2 hours

OpenStudy (anonymous):

it is simply cool, MILLIONS use it, just watch the intro vid, and you will be hooked

OpenStudy (anonymous):

all i remember about calculus is that they asked to find the rate of change between point A and point A, and when it wasn't 0 I was throughly confused :)

OpenStudy (anonymous):

I will defiantly check that site out, and will likely return here with questions. I will do my best to fan you when i'm allowed to. Thanks for your time and patience! See ya soon

OpenStudy (anonymous):

http://www.khanacademy.org/#Calculus